A Parallel plate capacitor having plate area A & plate separation D is joined to a battery of emf E & internal resistance R at t=0 consider a plane surface of area A/2 parallel to the plates & situated symmetrically between them. Find the displacement current through this surface as a function of time?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Kunal
 
Sol. E = Kq/x^2 , [from coulomb’s law] ∅ base E = EA = KqA/x^2 I base d = ϵ base 0 d∅E/dt = ϵ base 0 d/dt kqA/x^2 = ϵ base 0 KqA = d/dt x^-2 = ϵ base 0 * 1/4π ϵ base 0 *q *A * -2 *x^-3 * dx/dt = qAv/2πx^3  RegardsArun (askIITians forum expert)